OpenStudy (anonymous):

solve the following. Use interval notation when appropriate: x^2-7x=12

OpenStudy (anonymous):

and \[\sqrt{3x-2}=7\]

myininaya (myininaya):

\[\sqrt{3x-2}=7\] Square both sides to get rid of the square root on that one side \[3x-2=7^2\]

myininaya (myininaya):

3x-2=49 now just add 2 on both sides and then after that divide by 3 on both sides

OpenStudy (anonymous):

how about \[x ^{2}-7x+12=0\]

myininaya (myininaya):

\[12=3(4)=-3(-4) ; -7=-3+(-4)\] so this means you can factor it

OpenStudy (anonymous):

how do I go from the equation I had to what you have?

myininaya (myininaya):

\[(x-4)(x-3)=0\]

OpenStudy (anonymous):

thank you!